### v1

@@ 6.2. Effect of Prompt Templates (Q2)
In this section, we aim to investigate the impact of different prompt templates on ADD performance. We employ prompts 1 to 5 as described in Sec. 3, omitting the final sentence ”Answer fake or real.”. To ensure that the results predominantly reflect the interaction between the prompt templates and the LLM, we froze the audio encoder and solely fine-tune the LLM component using the LoRA technique, which corresponds to $\mathcal{A}LLM4ADD^{\star}$. EER (%) on ASVspoof2019 LA evaluation set across different settings are presented in Table 5.
@@ 6.3. Effect of LoRA Rank (Q3)
To evaluate the impact of different LoRA ranks, we explore the rank with values $\{1,4,16,64,256\}$. Following Sec. 6.2, we employ $\mathcal{A}LLM4ADD^{\star}$ for experiments. Performance is evaluated using the ASVspoof2019 LA evaluation set, with EER (%) measured across the $ASV@full$, $ASV@1/4$, $ASV@1/8$, and $ASV@1/16$ settings. Experimental results are depicted in Figure 2.
@@ 6.4. Effect of ALLM backbones (Q4)
To explore the impact of different ALLM backbones, we compare the performance of the Qwen-audio-chat and Qwen-audio-base backbone at $ASV@full$ and $ASV@1/16$ settings. Table 6 presents the performance on ASVspoof2019 LA evaluation set. $\mathcal{A}LLM4ADD^{\star}$ indicates that the audio encoder is frozen while $\mathcal{A}LLM4ADD^{\triangle}$ denotes that it is trainable. Further evaluation of more ALLMs is reserved for future work.
Table 6. Comparison of different ALLM backbones. We report EER (%), ACC (%), and AUC (%) on ASVspoof2019 LA evaluation set across ASV@full and ASV@1/16 settings. Base and Chat denote Qwen-audio-base and Qwen-audio-chat backbones, respectively.
@@
$\mathcal{A}LLM4ADD^{\star}$(Base) 1.78 98.33 99.65 4.19 97.84 99.10
$\mathcal{A}LLM4ADD^{\star}$(Chat) 1.30 99.26 99.88 3.13 98.11 99.43
$\mathcal{A}LLM4ADD^{\triangle}$(Base) 1.29 98.25 99.71 3.79 98.63 99.21
$\mathcal{A}LLM4ADD^{\triangle}$(Chat) 0.41 99.39 99.97 2.45 98.52 99.21
@@
From Table 6, we observe that the Qwen-audio-chat backbone consistently outperforms the Qwen-audio-base backbone under both the $ASV@full$ and $ASV@1/16$ settings, regardless of whether the audio encoder is trainable. For example, under $\mathcal{A}LLM4ADD^{\triangle}$ setting, the Qwen-audio-chat backbone achieves an EER of 0.41%, compared with 1.29% for the Qwen-audio-base backbone when trained on $ASV@full$.
@@
$\mathcal{A}LLM4ADD^{\star}$ 9.10 14.53 1.00 1.33
$\mathcal{A}LLM4ADD^{\triangle}$ $\mathbf{8.14}$ $\mathbf{11.05}$ $\mathbf{0.86}$ $\mathbf{1.28}$
@@ 6.5. Performance on other Fake Datasets (Q5)
In this section, we aim to explore the performance of $\mathcal{A}LLM4ADD$ on different types of ADD datasets. Specifically, we focus on the EmoFake (Zhao et al., 2024) and SceneFake (Yi et al., 2024) datasets. EmoFake involves modifying the emotional characteristics of speech while preserving other information. SceneFake, on the other hand, entails altering the acoustic scene of an utterance via speech enhancement techniques, without changing other aspects. We evaluate both $\mathcal{A}LLM4ADD^{\star}$ and $\mathcal{A}LLM4ADD^{\triangle}$, alongside several baselines under both @$full$ and @1/16@116@1/16@ 1 / 16 settings. For the @1/16@116@1/16@ 1 / 16 setting of EmoFake and SceneFake, we extract 1/16 of the real and fake audio files from the corresponding training set. Performance on the corresponding evaluation set is presented in Table 7.
@@
We observe that the $\mathcal{A}LLM4ADD^{\triangle}$ consistently achieves the lowest EER across both SceneFake and EmoFake datasets, under both @$full$ and @1/16@116@1/16@ 1 / 16 settings. Specifically, at the @$full$ setting, $\mathcal{A}LLM4ADD^{\triangle}$ exhibits an EER of 8.14% for SceneFake and 0.86% for EmoFake, compared to the best EERs of competing end-to-end models, which are 11.38% and 1.22%, respectively. This consistent superior performance underscores the effectiveness of our approach, validating its capability in detecting various types of audio deepfakes.
Figure 3. Performance on ASVspoof 2019 LA evaluation set across different data ratios. We depict EER (%) and AUC (%) of $\mathbf{\mathcal{A}LLM4ADD^{\triangle}}$ and $\mathbf{\mathcal{A}LLM4ADD^{\star}}$.
@@ 6.6. Performance on extremely limited data (Q6)
Although we report the performance of $\mathcal{A}LLM4ADD$ under various training set sizes in Sec. 5.3, the model’s effectiveness on extremely limited data volume is not explored. To address this, we conduct further experiments with training sets sized at fractions {1, 1/2, 1/4, 1/8, 1/16, 1/32, 1/64, 1/128} of the total data volume in this section. For $ASV@1/32$, $ASV@1/64$, and $ASV@1/128$ settings, we train the models for 15 epochs and adjust the learning rate according to the data volume to ensure the best performance. Notably, under the $ASV@1/128$ setting, our training dataset consists of only 178 fake audio samples and 20 real audio samples. Figure 3 illustrates the EER (%) and ACC (%) of $\mathcal{A}LLM4ADD^{\triangle}$ and $\mathcal{A}LLM4ADD^{\star}$ across these proportions.
@@
From Figure 3, we draw the following observations: (1) The performance of the model tends to decline as the data ratio decreases. (2) Our method still achieves impressive performance even in settings with extremely scarce data. For instance, in the $ASV@1/128$ setting, $\mathcal{A}LLM4ADD^{\triangle}$ still maintains an EER below 5% and an accuracy over 96%. This can likely be attributed to the ALLM’s robust few-shot capabilities, which enable it to adapt effectively to downstream tasks with little additional training. These experimental results demonstrate the feasibility of effective audio deepfake detection in scenarios with minimal data.
@@ 7. Conclusion
This paper presents our pioneering work on applying ALLMs to ADD. First, we conduct a comprehensive evaluation of ALLMs’ zero-shot capabilities for fake audio detection, revealing their limitations on the ADD task. We then propose $\mathcal{A}LLM4ADD$, a novel framework that reformulates the ADD task as an AQA problem, to endow ALLMs with the ability to detect fake audio. Extensive empirical results demonstrate that $\mathcal{A}LLM4ADD$ can achieve superior performance compared to existing methods, particularly in data-scarce scenarios. Notably, our method can achieve an EER below 5% and an accuracy over 96% with only around 200 training samples. These findings underscore the potential of ALLM-based approaches in advancing fake audio detection. In the future, we plan to leverage ALLMs to develop an ADD system that unifies deepfake detection with explaining capabilities.

### v2

@@ 6.2. Effect of Prompt Templates (Q2)
In this section, we aim to investigate the impact of different prompt templates on ADD performance. We employ prompts 1 to 5 as described in Sec. 3, omitting the final sentence ”Answer fake or real.”. To ensure that the results predominantly reflect the interaction between the prompt templates and the LLM, we froze the audio encoder and solely fine-tune the LLM component using the LoRA technique, which corresponds to $ALLM4ADD^{\star}$. EER (%) on ASVspoof2019 LA evaluation set across different settings are presented in Table 5.
@@ 6.3. Effect of LoRA Rank (Q3)
To evaluate the impact of different LoRA ranks, we explore the rank with values $\{1,4,16,64,256\}$. Following Sec. 6.2, we employ $ALLM4ADD^{\star}$ for experiments. Performance is evaluated using the ASVspoof2019 LA evaluation set, with EER (%) measured across the $ASV@full$, $ASV@1/4$, $ASV@1/8$, and $ASV@1/16$ settings. Experimental results are depicted in Figure 2.
@@ 6.4. Effect of ALLM backbones (Q4)
To explore the impact of different ALLM backbones, we compare the performance of the Qwen-audio-chat and Qwen-audio-base backbone at $ASV@full$ and $ASV@1/16$ settings. Table 6 presents the performance on ASVspoof2019 LA evaluation set. $ALLM4ADD^{\star}$ indicates that the audio encoder is frozen while $ALLM4ADD^{\triangle}$ denotes that it is trainable. Further evaluation of more ALLMs is reserved for future work.
Table 6. Comparison of different ALLM backbones. Base and Chat denote Qwen-audio-base and Qwen-audio-chat backbones, respectively.
@@
$ALLM4ADD^{\star}$(Base) 1.78 98.33 99.65 4.19 97.84 99.10
$ALLM4ADD^{\star}$(Chat) 1.30 99.26 99.88 3.13 98.11 99.43
$ALLM4ADD^{\triangle}$(Base) 1.29 98.25 99.71 3.79 98.63 99.21
$ALLM4ADD^{\triangle}$(Chat) 0.41 99.39 99.97 2.45 98.52 99.21
@@
From Table 6, we observe that the Qwen-audio-chat backbone consistently outperforms the Qwen-audio-base backbone under both the $ASV@full$ and $ASV@1/16$ settings, regardless of whether the audio encoder is trainable. For example, under $ALLM4ADD^{\triangle}$ setting, the Qwen-audio-chat backbone achieves an EER of 0.41%, compared with 1.29% for the Qwen-audio-base backbone when trained on $ASV@full$.
@@
$ALLM4ADD^{\star}$ 9.10 14.53 1.00 1.33
$ALLM4ADD^{\triangle}$ $\mathbf{8.14}$ $\mathbf{11.05}$ $\mathbf{0.86}$ $\mathbf{1.28}$
@@ 6.5. Performance on other Fake Datasets (Q5)
In this section, we aim to explore the performance of ALLM4ADD on different types of ADD datasets. Specifically, we focus on the EmoFake (Zhao et al., 2024) and SceneFake (Yi et al., 2024) datasets. EmoFake involves modifying the emotional characteristics of speech while preserving other information. SceneFake, on the other hand, entails altering the acoustic scene of an utterance via speech enhancement techniques, without changing other aspects. We evaluate both $ALLM4ADD^{\star}$ and $ALLM4ADD^{\triangle}$, alongside several baselines under both @$full$ and @1/16@116@1/16@ 1 / 16 settings. For the @1/16@116@1/16@ 1 / 16 setting of EmoFake and SceneFake, we extract 1/16 of the real and fake audio files from the corresponding training set. Performance on the corresponding evaluation set is presented in Table 7.
@@
We observe that the $ALLM4ADD^{\triangle}$ consistently achieves the lowest EER across both SceneFake and EmoFake datasets, under both @$full$ and @1/16@116@1/16@ 1 / 16 settings. Specifically, at the @$full$ setting, $ALLM4ADD^{\triangle}$ exhibits an EER of 8.14% for SceneFake and 0.86% for EmoFake, compared to the best EERs of competing end-to-end models, which are 11.38% and 1.22%, respectively. This consistent superior performance underscores the effectiveness of our approach, validating its capability in detecting various types of audio deepfakes.
Figure 3. Performance on ASVspoof 2019 LA evaluation set across different data ratios. We depict EER (%) and AUC (%) of $\mathbf{ALLM4ADD^{\triangle}}$ and $\mathbf{ALLM4ADD^{\star}}$.
@@ 6.6. Performance on extremely limited data (Q6)
Although we report the performance of ALLM4ADD under various training set sizes in Sec. 5.3, the model’s effectiveness on extremely limited data volume is not explored. To address this, we conduct further experiments with training sets sized at fractions {1, 1/2, 1/4, 1/8, 1/16, 1/32, 1/64, 1/128} of the total data volume in this section. For $ASV@1/32$, $ASV@1/64$, and $ASV@1/128$ settings, we train the models for 15 epochs and adjust the learning rate according to the data volume to ensure the best performance. Notably, under the $ASV@1/128$ setting, our training dataset consists of only 178 fake audio samples and 20 real audio samples. Figure 3 illustrates the EER (%) and ACC (%) of $ALLM4ADD^{\triangle}$ and $ALLM4ADD^{\star}$ across these proportions.
@@
From Figure 3, we draw the following observations: (1) The performance of the model tends to decline as the data ratio decreases. (2) Our method still achieves impressive performance even in settings with extremely scarce data. For instance, in the $ASV@1/128$ setting, $ALLM4ADD^{\triangle}$ still maintains an EER below 5% and an accuracy over 96%. This can likely be attributed to the ALLM’s robust few-shot capabilities, which enable it to adapt effectively to downstream tasks with little additional training. These experimental results demonstrate the feasibility of effective audio deepfake detection in scenarios with minimal data.
@@ 7. Conclusion
This paper presents our pioneering work on applying ALLMs to ADD. First, we conduct a comprehensive evaluation of ALLMs’ zero-shot capabilities for fake audio detection, revealing their limitations on the ADD task. We then propose ALLM4ADD, a novel framework that reformulates the ADD task as an AQA problem, to endow ALLMs with the ability to detect fake audio. Extensive empirical results demonstrate that ALLM4ADD can achieve superior performance compared to existing methods, particularly in data-scarce scenarios. Notably, our method can achieve an EER below 5% and an accuracy over 96% with only around 200 training samples. These findings underscore the potential of ALLM-based approaches in advancing fake audio detection. In the future, we plan to leverage ALLMs to develop an ADD system that unifies deepfake detection with explaining capabilities.
8. Acknowledgements
This work is supported by the National Natural Science Foundation of China (NSFC) (No. 62322120, No.U21B2010, No. 62306316, No. 62206278).